OpenStudy (anonymous):

the measure of an angle is 34 greater than its complement find the measure of each angle

OpenStudy (asnaseer):

do you know what complementary angles are?

OpenStudy (anonymous):

yes. 90 degree angle

OpenStudy (asnaseer):

sort of - two angles are complementary if they add up to 90 degrees

OpenStudy (anonymous):

let the two angels are "X" and "Y" and "X">"Y" then; X-Y=34 ; and X+Y=90 by adding the two equations then 2X=124 then X=62 and y=28

OpenStudy (asnaseer):

so if you call one of the angles x

OpenStudy (asnaseer):

then your question can be formulated as: x = 34 + (90-x)

OpenStudy (asnaseer):

does that make sense?

OpenStudy (asnaseer):

this is because if one angle is x, then its complement must be (90-x)

OpenStudy (anonymous):

YES!!! thank you guys!! you both are awesome!!

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

yw bro

OpenStudy (anonymous):

you will be assigned 3 projects in your literature class this semester.If you earn a 92% and 84% on the first two, what will you have to earn on the third to earn at least a 90% project average? can you guys help me with this problem?!

OpenStudy (asnaseer):

please post each question separately

OpenStudy (asnaseer):

in the list on the left

OpenStudy (anonymous):

why?

OpenStudy (asnaseer):

the code of conduct says you should post one question at a time. it also helps others on this site to see your individual questions and gain insights from them.

OpenStudy (anonymous):

ohh ok. I'm sorry. this is my first time on this website.. so yea.

OpenStudy (asnaseer):

thank you. :)

OpenStudy (asnaseer):

since you are fairly new you also ought to know that the usual workflow is: 1. You post a question. 2. One or more people come to your question to help you out. 3. You select the one that best answered your question by clicking on the blue Best Response button next to any one of their replies. 4. You close the question by clicking on the Close button then the whole process starts again. In addition, if you are helping others on this site, you should try and guide them to an answer rather than work everything out for them. That way the asker learns more. And lastly - Welcome to OpenStudy! :)

OpenStudy (anonymous):

Thank you! :)

OpenStudy (asnaseer):

yw :)